OpenStudy (anonymous):

*What number shows up most often when you roll 10 dice?

OpenStudy (anonymous):

Do you mean the total of the ten? Average 1 die: (1 + 2 + 3 + 4 +5 + 6)/6 = 21/6 = 3.5, so 10 x 3.5 = 35 seems most likely. Also, 7 most likely for two dice, and 5 x 7 = 35.

OpenStudy (anonymous):

thank you <3